# Channel-wise Parallelizable Spiking Neuron with Multiplication-free Dynamics and Large Temporal Receptive Fields

### v1

@@ 4.2 Multiplication-Free Neuronal Dynamics
One of the most attractive characteristics of SNNs is that the multiply-accumulate (MAC) operations between binary spikes and synaptic weights can be superseded by accumulate (AC) operations during inference in neuromorphic chips (Pei et al., 2019). However, computational costs of the neuronal dynamics have not been paid much attention. For example, the PSN, the sliding PSN, and the Stochastic PSN involve massive MAC operations between floating-point neuronal weights and input currents. To future reduce the internal computation costs of spiking neurons, we introduce the bit-shift operation to supersede multiplication, which has been successfully employed in quantized neural networks (You et al., 2020; Elhoushi et al., 2021).
Figure 3: The workflow of power-of-2 quantizer.
When $w$ is the power of 2, then the multiplication of $w$ and $x$ can be replaced by the bit shift operation:
@@
where <<much-less-than$<$ < is the left bit shift operation. In particular, when $\log_{2}(w)<0$, left shifting an negative number $\log_{2}(w)$ of bits is actually right shifting $|\log_{2}(w)|$. To employ the bit-shift operation, we quantize $\bm{W}$ in Eq.(10) to $\bm{W}_{q}$, whose elements are the power of $2$, by an quantizing function $\mathcal{Q}$:
@@
where $\mathrm{sign}(x)$ is the sign function and returns the sign ($1$ or $-1$) of the input $x$; $\mathrm{round}(x)$ is the rounding-to-nearest function. Figure 3 shows the workflow of Eq.(13).
Figure 4: Using the Straight-Through Estimator to redefine $\mathrm{round}^{\prime}(x)$ solely will still cause a discrete gradient $\mathcal{Q}^{\prime}(x)$.
Remarkably, the gradient of $\mathrm{round}(x)$ is zero almost everywhere, and other operations in Eq.(13) are differentiable. The standard practice is to employ the Straight-Through Estimator (Bengio et al., 2013) to redefine its gradient as 1:
@@
However, Eq.(15) is still unstable because $\mathrm{round}(x)$ causes jump points and it oscillates around 0, shown in Figure 4.
To avoid the numerical instability caused by Eq.(15), we redefine $\mathcal{Q}^{\prime}(x)$ as a whole Straight-Through Estimator, rather than using Eq.(14) solely:
@@
where $\bm{V}_{th}\in\mathbb{R}^{C}$ is the learnable channel-wise threshold, and $\log_{2}(\bm{W}_{q})\in\mathbb{Z}^{C\times k}$ is quantized from $\bm{W}$ by Eq.(13). Note that $\log_{2}(\bm{W}_{q})$ is solved beforehand and there are no $\log$ and multiplication operations during inference.
@@ 4.3 Training Acceleration
The motivation for proposing parallelizable spiking neuron models is to solve the slow training speed of SNNs on GPUs caused by the step-by-step iterations of traditional serial neuron models. Thus, the efficient implementation of mul-free channel-wise PSN requires elaborate consideration.
@@
The computations of mul-free channel-wise PSN focus on Eq.(17), which is a typical 1-D convolution. The straightforward implementation uses PyTorch’s 1-D convolution (Conv1d) for the proposed neuron layer and the prevailing spiking deep learning framework SpikingJelly (Fang et al., 2023a) for other modules. However, Conv1d requires the sequence length, which is the time dimension, as the last dimension. While the default data layout in SpikingJelly is the time-first layout, using $(T,N,...)$ to represent the sequence data, where $T$ is the length, $N$ is the batch size, and ”...” represents any additional dimensions. This layout is necessary for accelerating stateless layers by fusing time and batch dimensions for parallel computing (Fang et al., 2023a). In this circumstance, the straightforward implementation is:
(1) Time-first + Reshape + Conv1d: this method uses PyTorch’s Conv1d, which requires the time-step at the last dimension. Thus, the reshape operations $(T,N,C,...)\rightleftharpoons(N*,C,T)$ before and after Conv1d are unavoidable, where $N*$ represents any additional dimensions in ”...” have been fused to the $N$ dimension into $N*$.
Unfortunately, the reshape operations in implementation (1) for the proposed neuron involve nonadjacent dimensions, resulting in slow memory reading/writing (R/W). Note that physical memory is 1-D, resulting in the fact that data in nonadjacent dimensions are also nonadjacent in physical memory. This is also the reason why SpikingJelly requires the time-first layout: the time and batch dimension fusion requires reshape operations $(T,N,...)\rightleftharpoons(TN,...)$, which involves adjacent dimensions and gets rid of memory R/W. Figure 5 demonstrates the cases of reshape operations with or without memory R/W.
Different implementations require different memory layouts, and consequentially lead to speed difference, even if the theoretical FLOPs are identical. Thus, the choice of data layouts is the primary principle. Different from the time-first layout in SpikingJelly, the time-last layout uses $(N,...,T)$ to represent sequences, which can also be considered. Based on the aforementioned background, we design the following five more implementations as candidates:
@@
(2) Time-first + Reshape + Conv2d: this method uses PyTorch’s 2-D convolution (Conv2d). It involves the reshape operations $(T,N,C,...)\rightleftharpoons(N,C,*,T)$ or $(T,N,C,...)\rightleftharpoons(N,C,T,*)$, and sets weight and stride in the ”*” dimension as 1. Compared with the implementation (1), its memory copying cost is less, while the 2-D convolution is more costly.
(3) Time-last + Vmap + Conv1d: this method uses the vectorizing map function (Vmap) in PyTorch to vectorize Conv1d to process the input sequence with the $(N,C,*,T)$ layout over the ”*” dimension. The reshape operations $(N,C,...,T)\rightleftharpoons(N,C,*,T)$ are nearly cost-free because the reshaped dimensions are adjacent physically.
(4) Time-last + Conv2d: this method is similar to the implementation (3), but processes the input sequence with the $(N,C,*,T)$ by Conv2d and sets the weight and stride in the ”*” dimension as 1, rather than by Vmap.
(5) Time-first/last + Vmap + MM: this method uses Vmap to vectorize matrix multiplication (MM) to process inputs over channels ($c$ in Eq.(17)). Refer to the Appendix for more details about how the weights for MM are generated.
(6) Time-first/last + Custom CUDA Kernel: this method avoids reshape operations and can be used for any memory layout. However, the convolutions in PyTorch are highly optimized, i.e. implemented by the official NVIDIA CUDA Deep Neural Network (cuDNN) library, which might be much faster than custom implementations.
Note that the implementations (3)-(6) adopt the time-last layout. Then the stateless layers should also use the same layout. Otherwise, reshape operations between time-first and time-last layouts will cause great latency. Correspondingly, the time batch dimension fusion method to accelerate stateless layers in SpikingJelly cannot be applied. We also elaborate on acceleration methods for stateless layers in the time-last layout, and the details can be found in the Appendix. As the acceleration of mul-free channel-wise PSN on GPUs will influence the stateless layers mutually, and the acceleration effect varies with the input shapes and GPUs, choosing of acceleration methods is empirical. We design an autoselect acceleration algorithm to avoid manually choosing, as shown in Algorithm 1. When the training of an SNN starts, the shape of the input sequence is also determined. Then this algorithm will run a benchmark over layouts and acceleration methods and select the options with the fastest speed.
[…]
Hammouamri et al. Two-layer FC + LIF + Learned Delay 95.10
Li et al. (2024b) Four-layer FC + RPSU 92.49
Chen et al. (2024b) Two-layer FC + PMSN 95.10
Yarga and Wood (2023) Two-layer FC + Stochastic PSN + Learned Delay 95.01
Ours Two-layer FC + Mul-free Channel-wise PSN + Learned Delay 95.71
@@
Datasets Mul-free Channel-wise PSN PMSN PSN masked PSN sliding PSN GLIF PLIF LIF
@@
Tan et al. (2022) ResNet-18 (ANN) BiGRU (ANN) 72.1
Bulzomi et al. (2023) Modified Spiking ResNet + PLIF FC (Stateful Synapses) 60.2
Dampfhoffer et al. (2024) ResNet-18 (ANN) BiGRU (ANN) 75.1
@@
Spiking ResNet-18 + PLIF SpikGRU2+ (Bi-direction + Sigmoid Gates + Ternary Spikes) 75.3
Ours Modified Spiking ResNet-18 + Mul-free Channel-wise PSN FC (Stateful Synapses) 70.9
5 Experiments
In this section, we evaluate the mul-free channel-wise PSN on various kinds of datasets. We conduct the ablation experiments on the order $k$ and demonstrate that the sawtooth dilations can compensate the long-term dependencies learning ability. Finally, we provide a training speed comparison to validate the efficiency of the autoselect algorithm.
5.1 Learning Long-Term Dependencies
We evaluate the long-term dependencies learning ability of mul-free channel-wise PSN in three widely used classification tasks, including the Spiking Heidelberg Digits (SHD) spoken digit dataset (Cramer et al., 2022), the sequential CIFAR dataset, and the high temporal resolution automatic lip-reading DVS-Lip dataset (Tan et al., 2022). These datasets cover the types of voices, images, and neuromorphic events.
Comparison between our methods and previous SOTA SNN methods on the SHD dataset are shown in Table 1. Specifically, we replace the LIF neurons in the SNN-delay architecture (Hammouamri et al., ) with our neurons. We achieve test accuracies of 95.31%, 95.62%, and 95.71% with sawtooth dilations and order $k=2,4,8$, respectively. To the best of our knowledge, these results represent the SOTA performance of SNN models on the SHD datasets.
Sequential image classification tasks have been commonly benchmarks to evaluate spiking neurons by (Yin et al., 2021; Fang et al., 2023b; Chen et al., 2024b). In these tasks, images are fed into the model column by column, and the number of time-step is equal to the width of the images. We also conduct experiments on sequential CIFAR10 and CIFAR100 datasets. To ensure fairness, we fully employ the network architecture and hyperparameters as (Fang et al., 2023b), only replacing the spiking neurons with ours. The results are shown in Table 2, where the data for PMSN is sourced from (Chen et al., 2024b), maintaining the same architecture as well, while the data for other neurons is sourced from (Fang et al., 2023b). On the sequential CIFAR10 dataset, our mul-free channel-wise PSN outperforms PSN by $2.72\%$ and PMSN by $0.2\%$. Additionally, on the sequential CIFAR100 dataset, it surpasses PSN by $4\%$ and PMSN by $0.13\%$. Notably, the order of our neurons here we report is 16, while the order of sliding PSN and masked PSN is 32, 2×2$\times$2 × than us.
Furthermore, we demonstrate the capability of mul-free channel-wise PSN in processing complex neuromorphic DVS-Lip dataset, which comprises 100 classes and consists of 19871 samples, each containing approximately $10^{4}$ events. These events are generated with a spatial resolution of $128\times 128$ pixels and a temporal resolution at the microsecond level. Half of the words in the dataset are visually similar pairs in the LRW dataset (Chung and Zisserman, 2017) (e.g., ”America” and ”American”). The training and testing sets are derived from different speakers, posing a challenge for the model to exhibit robust generalization capabilities with respect to speaker characteristics.
Currently, the SOTA accuracy of 75.3% on the DVS-Lip dataset is achieved by (Dampfhoffer et al., 2024) using a Spiking ResNet-18 with the channel-wise PLIF neurons fronted, a SpikGRU2+ backend, and events are integrated into 90 frames ($T=90$). In our experiments, we introduce several modifications to the frontend. We replace the PLIF neurons with our neurons and remove spiking neurons from the pooling layers, referring to this architecture as Modified Spiking ResNet-18. As Table 3 shows, our method achieves 70.9% accuracy and is only second to (Dampfhoffer et al., 2024) with SpikGRU2+ backend. It is worth noting that SpikGRU2+ is bi-directional with two groups of separate hidden states, employs sigmoid gates with floating activations, and outputs ternary spikes ($-1,1,0$), which is not a pure SNN module and might be difficult to be deployed to neuromorphic chips. The accuracy we report here is based on the neuron order $k=2$ and sawtooth dilations, indicating that our method can effectively capture rich historical information with a small order even when handling tasks involving long-time sequences.
@@ 5.2 Ablation Study
To evaluate the effectiveness of sawtooth dilations, we conduct an ablation study on the pixel CIFAR10 classification task. In this task, images are flattened into one-dimensional vectors as time series inputs to the network. Thus, the number of time-step is $T=1024$. We adopted the same network structure as (Chen et al., 2024b). Figure 6 (b) illustrates the accuracy curves with/without sawtooth dilations. It can be observed that the accuracy with sawtooth dilations is consistently higher than that without. Additionally, when the order $k$ is small, which is a practical case for deployment, the accuracy of our neuron with sawtooth dilations is much higher. These results validate that the sawtooth dilations compensate the effect of large receptive fields when using a small $k$.
@@ 5.3 Training Acceleration
We compare the training speed of PSN and the mul-free channel-wise PSN implemented by the autoselect Algorithm 1. The naive manual implementations, using reshape operations and PyTorch’s 1-D convolutions for neuron layers and time batch dimension fusion or the vectorizing map for stateless layers, are also compared. The experiments are carried out on a Debian GNU/Linux 11(bullseye) server with an Intel(R) Xeon(R) Platinum 8336C CPU, a Nvidia A100-SXM4-80GB GPU and 32GB RAM. We set the batch size as 128.
@@
The training duration ($s$/epoch) of different neurons under different order $k$ on CIFAR100 is shown in Figure 7. Note that both PSN and sliding PSN are implemented by matrix multiplication in GPUs (Fang et al., 2023b), their speeds are identical and decoupled with $k$. For the sake of easy reading, we plot the neurons with the same $k$ in similar styles. The results show that our autoselect algorithm greatly improves the efficiency of mul-free channel-wise PSN and achieves a much faster training speed than the naive manual implementations. When $T\leq 4$, our method is comparable to PSN/Sliding PSN. In this case, the matrixs are nearly strips in PSN/Sliding PSN because $T$ is much less than other dimensions, causing inefficient matrix multiplications. When $T$ continuously increases, our method is slower, which is caused by the fact that the quantization induces some additional overhead, and memory reading/writing caused by reshape or vectorizing map operations for processing inputs/outputs in our SNNs is slower than the dimension fusion. Nonetheless, the speed gaps are not significant. The high task accuracy and hardware-friendly advantages make mul-free channel-wise PSN a strong alternative for PSN and sliding PSN.

### v2

@@ 4.2 Multiplication-Free Neuronal Dynamics
(a) The workflow of power-of-2 quantizer.
(b) Using Straight-Through Estimator (STE) to redefine $\mathrm{round}^{\prime}(x)$ will cause a discrete gradient $\mathcal{Q}^{\prime}(x)$.
Figure 3: Power-of-2 quantization and the gradient behavior of $\mathrm{round}^{\prime}(x)$ under STE.
To future reduce the internal computation costs of spiking neurons, we introduce the bit-shift operation to supersede multiplication, which has been successfully employed in quantized neural networks [33, 34]. In this way, the entire network can perform inference without any multiplication operations. For multiplication of IEEE 754-defined FP32/FP16 values $x$ with powers of 2 (denoted as $w$), the operation can be converted to a lower-bit integer addition to the floating-point exponent bits [35]. For integers $x$ multiplied by $w$, it can be achieved by the bit-shift operation, as shown in Eq.(12).
@@
where <<<< is the left bit-shift operation. In particular, when $\log_{2}(w)<0$, left shifting an negative number $\log_{2}(w)$ of bits is actually right shifting $|\log_{2}(w)|$. To employ the bit-shift operation, we quantize $\bm{W}$ in Eq.(10) to $\bm{W}_{q}$, whose elements are the power of $2$, by an quantizing function $\mathcal{Q}$:
@@
where $\mathrm{sign}(x)$ is the sign function and returns the sign ($1$ or 1-1) of the input $x$; $\mathrm{round}(x)$ is the rounding-to-nearest function. Figure 3(a) shows the workflow of Eq.(13).
Remarkably, the gradient of $\mathrm{round}(x)$ is zero almost everywhere, and other operations in Eq.(13) are differentiable. The standard practice is to employ the Straight-Through Estimator [36] to redefine its gradient as 1:
@@
However, Eq.(15) is still unstable because $\mathrm{round}(x)$ causes jump points and it oscillates around 0, shown in Figure 3(b), which may cause the collapse. To avoid the numerical instability caused by Eq.(15), we redefine $\mathcal{Q}^{\prime}(x)$ as a whole Straight-Through Estimator, rather than using Eq.(14) solely:
@@
where $\bm{V}_{th}\in\mathbb{R}^{C}$ is the learnable channel-wise threshold, and $\log_{2}(\bm{W}_{q})\in\mathbb{Z}^{C\times k}$ is quantized from $\bm{W}$ by Eq.(13). Note that $\log_{2}(\bm{W}_{q})$ is solved beforehand and there are no $\log$ and multiplication operations during inference. Table 1 presents an analysis of the parameter and operation costs of various neurons, evaluated under $T$ time-steps, $k$ order and $C$ channels.
Table 1: Parameters and computational costs of different spiking neurons during inference.
Neuron Params Operations
PSN $T^{2}+T$ $(T^{2}+T)\times\text{ADD},T^{2}\times\text{MUL}$
Sliding PSN $k+1$ $((T+\frac{1-k}{2})\cdot k+T)\times\text{ADD},((T+\frac{1-k}{2})\cdot k)\times\text{MUL}$
Ours $C\cdot(k+1)$ $((T+\frac{1-k}{2})\cdot k+T)\times\text{ADD},((T+\frac{1-k}{2})\cdot k)\times\text{SHIFT}$
@@ 4.3 Training Acceleration
The motivation for proposing parallelizable spiking neuron models is to solve the slow training speed of SNNs on GPUs caused by the step-by-step iterations of traditional serial neuron models. Efficient implementation of mul-free channel-wise PSN, a typical 1-D convolution described in Eq.(17), requires elaborate consideration. To tackle this, we explore the implementations under two SNN data layouts: time-first and time-last. The time-first layout (T,N,C,)(T,N,C,...), widely adopted in SNN frameworks like SpikingJelly [6], accelerates stateless layers by fusing the $T$ (time-steps) and $N$ (batch size) dimensions, while $C$ denotes the channel dimension and "..." represents any additional dimensions. In comparison, the time-last layout (N,C,,T)(N,C,...,T) arranges $T$ as the last dimension.
A vanilla implementation of Eq.(17) leverages PyTorch’s 1-D Convolution (Conv1d), which requires the shape of inputs as $(N,C,T)$. However, for both time-first and time-last layout, it is unavoidable to perform reshape operations $(T,N,C,...)\rightleftharpoons(N*,C,T)$ and $(N,C,...,T)\rightleftharpoons(N*,C,T)$ before and after the Conv1d to satisfy the shape requirement of Conv1d, where NN* represents the fusion of any additional dimensions in "..." into the $N$ dimension. Note that physical memory is 1-D, meaning that data in nonadjacent dimensions is also stored nonadjacent in physical memory. These reshape operations involve nonadjacent dimensions and require costly memory reading/writing (R/W) operations to create a new contiguous tensor in physical memory. Figure 4 illustrates examples of reshape operations with or without memory R/W.
@@
To address the overhead caused by costly nonadjacent reshape operations in the neuron layer, we have developed several acceleration methods. For the time-first layout (T,N,C,)(T,N,C,...), we propose two methods. One is a custom CUDA kernel that directly performs convolutions along the $T$ dimension. Another leverages PyTorch’s vectorizing map function (Vmap) to parallelize computations over the $C$ dimension, followed by matrix multiplication (MM) to process other dimensions. For the time-last layout (N,C,,T)(N,C,...,T), in addition to the custom CUDA kernel or Vamp + MM approach as in the time-first layout, we propose two additional methods. One uses the 2-D convolution to implement the 1-D convolution with the weight and stride as 1 to handle "..." dimension. Another applies Vmap to vectorize the $C$ dimension and employs Conv1d to handle other dimensions. Furthermore, we also elaborate on acceleration strategies for stateless layers in the time-last layout. Details for accelerating neuron and stateless layers are provided in Appendix A and B, respectively.
The choice of acceleration methods for neuron layers affects the data layout, which subsequently influences the performance of stateless layers. Moreover, the acceleration effect is dependent on input shapes and GPUs, making the process of selecting acceleration methods inherently empirical. To eliminate the need for manual selection, we design an autoselect acceleration algorithm, as shown in Algorithm 1. At the start of SNN training, the input sequence’s shape is determined. Then the algorithm will run a benchmark over data layouts and acceleration methods, selecting the configuration with the highest execution speed. In Algorithm 1, each $t_{l,i}$ is evaluated through $2m+1$ repeated executions, with the average execution time of the last $m$ iterations serving as the measured runtime for the corresponding acceleration method. As the autoselect method is invoked only once during the entire training phase, its impact on overall train time is negligible.
@@
5 Experiments
In this section, we evaluate the mul-free channel-wise PSN on various kinds of datasets. We conduct the ablation experiments on the order $k$ and demonstrate that the sawtooth dilations can compensate for the long-term dependencies learning ability. Finally, we provide a training speed comparison to validate the efficiency of the autoselect algorithm.
5.1 Learning Long-Term Dependencies
We evaluate the long-term dependencies learning ability of mul-free channel-wise PSN in three widely used classification tasks, including the Spiking Heidelberg Digits (SHD) spoken digit dataset [37], the sequential CIFAR dataset, and the high spatial-temporal resolution automatic lip-reading DVS-Lip dataset [38]. These datasets cover the types of voices, images, and neuromorphic events.
Comparison between our method and previous SOTA SNN methods on the SHD dataset is shown in Table 2. Specifically, we replace the LIF neurons in the SNN-delay architecture [39] with our neurons. With sawtooth dilations and order $k=2$, we achieve a test accuracy of $95.50\pm 0.36\%$ under three seeds ($0,1,2$).
Sequential image classification tasks have been commonly benchmarks to evaluate spiking neurons by [40, 15, 29]. In these tasks, images are fed into the model column by column, and the number of time-step is equal to the width of the images. We also conduct experiments on sequential CIFAR10 and CIFAR100 datasets. To ensure fairness, we fully employ the network architecture and hyperparameters as [15], only replacing the spiking neurons with ours. The results are shown in Table 3, where the data for PMSN is sourced from [29], maintaining the same architecture as well, while the data for other neurons is sourced from [15]. On the sequential CIFAR10 dataset, our mul-free channel-wise PSN outperforms PSN by $2.72\%$ and PMSN by $0.2\%$. Additionally, on the sequential CIFAR100 dataset, it surpasses PSN by $4\%$ and PMSN by $0.13\%$. Notably, the order of our neurons here we report is 16, while the order of sliding PSN and masked PSN is 32, $2\times$ than ours.
Furthermore, we demonstrate the capability of mul-free channel-wise PSN in processing complex neuromorphic DVS-Lip dataset, which comprises 100 classes and consists of 19871 samples, each containing approximately $10^{4}$ events with a spatial resolution of $128\times 128$. Half of the words in the dataset are visually similar pairs in the LRW dataset [41] (e.g., "America" and "American"). The training and testing sets are derived from different speakers, posing a challenge for the model to exhibit robust generalization capabilities with respect to speaker characteristics.
Currently, the SOTA accuracy of 75.3% on the DVS-Lip dataset is achieved by [42] using a Spiking ResNet-18 with the channel-wise PLIF neurons fronted, a SpikGRU2+ backend, and events are integrated into 90 frames ($T=90$). In our experiments, we introduce several modifications to the frontend. We replace the PLIF neurons with our neurons and remove spiking neurons from the pooling layers, referring to this architecture as Modified Spiking ResNet-18. As Table 4 shows, our method achieves 70.9% accuracy and is only second to [42] with SpikGRU2+ backend. It is worth noting that SpikGRU2+ is bi-directional with two groups of separate hidden states, employs sigmoid gates with floating activations, and outputs ternary spikes (1,1,0-1,1,0), which is not a pure SNN module and might be difficult to deploy to neuromorphic chips. The accuracy we report here is based on the neuron order $k=2$ and sawtooth dilations, indicating that our method can effectively capture rich historical information with a small order even when handling tasks involving long-time sequences.
@@ 5.1 Learning Long-Term Dependencies
Method Network Parallel Accuracy($\%$)
Hammouamri et al. [39] Two-layer FC + LIF + Learned Delay $95.07\pm 0.24$
Li et al. [30] Four-layer FC + RPSU $92.49$
Chen et al. [29] Two-layer FC + PMSN $95.10$
Yarga and Wood [16] Two-layer FC + Stochastic PSN + Learned Delay $95.01$
Ours Two-layer FC + Mul-free Channel-wise PSN + Learned Delay $95.50\pm 0.36$
Datasets Ours PMSN[29] PSN[15] Masked PSN[15] Sliding PSN[15] GLIF[13] PLIF[5] LIF
@@
Tan et al. [38] ResNet-18 (ANN) BiGRU (ANN) 72.1
Bulzomi et al. [43] Modified Spiking ResNet + PLIF FC (Stateful Synapses) 60.2
Dampfhoffer et al. [42] ResNet-18 (ANN) BiGRU (ANN) 75.1
@@
Spiking ResNet-18 + PLIF
SpikGRU2+
(Bi-direction + Sigmoid Gates + Ternary Spikes)
75.3
Ours
Modified Spiking ResNet-18
+ Mul-free Channel-wise PSN
FC (Stateful Synapses) 70.9
@@ 5.2 Ablation Study
To evaluate the effectiveness of sawtooth dilations, we conduct an ablation study on the pixel CIFAR10 classification task. In this task, images are flattened into one-dimensional vectors as time series inputs to the network. Thus, the number of time-step is $T=1024$. We adopt the same network structure as [29]. Figure 5 (b) illustrates the accuracy curves with/without sawtooth dilations. It can be observed that the accuracy with sawtooth dilations is consistently higher than that without. Additionally, when the order $k$ is small, which is a practical case for deployment, the accuracy of our neuron with sawtooth dilations is much higher. These results validate that the sawtooth dilations compensate for the effect of large receptive fields when using a small $k$.
@@ 5.3 Training Acceleration
We compare the training speed of PSN and the mul-free channel-wise PSN implemented by the autoselect Algorithm 1. The vanilla implementation, using reshape and Conv1d operations for neuron layers and time batch dimension fusion for stateless layers, is also compared. The experiments are carried out on a Debian GNU/Linux 11(bullseye) server with an Intel(R) Xeon(R) Platinum 8336C CPU, an Nvidia A100-SXM4-80GB GPU and 32GB RAM. Following the original PSN [15] settings, we use the batch size of 128 on the CIFAR dataset.
@@
The training duration (ss/epoch) of different neurons under different order $k$ on CIFAR100 is shown in Figure 6. Note that both PSN and sliding PSN are implemented by matrix multiplication in GPUs [15], their speeds are identical and decoupled with $k$. The results show that our autoselect algorithm greatly improves the efficiency of mul-free channel-wise PSN and achieves a much faster training speed than the vanilla implementation. When $T\leq 4$, our method is comparable to PSN/sliding PSN. In this case, the matrix is nearly stripped in PSN/sliding PSN because $T$ is much less than other dimensions, causing inefficient matrix multiplications. When $T$ continuously increases, our method is slower, which is caused by the fact that the quantization induces additional overhead, and memory R/W caused by Vmap operations for processing inputs/outputs in our SNNs is slower than the dimension fusion. Nonetheless, the speed gaps are not significant. The autoselect algorithm itself introduces only negligible time overhead, contributing $1.17\%,0.39\%,0.27\%$ for $T=2,8,32$, respectively. Further results for different batch sizes and GPUs can be found in Appendix G.
5.4 Inference Memory Analysis
Both our approach and sliding PSN require storing an input sequence whose length is proportional to the neuron order $k$. According to [15], sliding PSN typically needs a large $k$ to maintain stable performance on long-term dependencies. In contrast, our method can achieve more stable and better performance with a much smaller $k$, as shown in Fig. 5. By substantially reducing the required neuron order, our approach shortens the length of the stored input sequence and significantly lowers memory consumption during inference. As demonstrated in Table 5, this leads to a substantial reduction in memory overhead for storing input sequences, making our solution more suitable for deployment on resource-constrained devices.
5.5 Inference Energy Estimation
Assume implemented on the 45 nm CMOS technology, where a 32-bit floating-point (FP32) multiplication (MUL) and addition (ADD) operation consumes $3.7$ pJ and $0.9$ pJ, respectively [44]. In comparison, for the shift operation of a 32-bit fixed-point (FIX32) and a power-of-two number, it requires only $0.13$ pJ [44]. For the multiplication of a FP32 number with a power-of-two number, it can be performed by one single lower-bit integer addition, which is also energy efficient [35]. Here we use the FIX32 shift energy as an approximation. Based on the sequential CIFAR100 models (see Appendix D), we estimate the average computational energy of PSN and our method for processing a single CIFAR100 image, as shown in Table 6. Our method achieves nearly $9\times$ lower energy consumption compared to PSN, demonstrating a significant advantage in hardware efficiency. Notably, our energy estimation considers the cost of neuronal layers, whereas most previous studies only account for synaptic layers. Moreover, we observe that for PSN, the energy consumption of the neuronal layer is significantly higher than that of the synaptic layer, which is mainly due to the dense floating-point matrix multiplications involved. Details of the energy estimation procedure can be found in Appendix F.
Table 5: Step-by-step inference memory on the sequential CIFAR100 ($T=32$) dataset.
Neuron $k$ Accuracy(%) Memory(MB)
Sliding PSN 32 62.11 2635
Ours 4 65.77 547
Table 6: Computational energy comparison on a single CIFAR100 Image.
Neuron Layer Synaptic Layer
Total Energy
[…]
PSN
$1.91\times 10^{7}$ MUL
$1.97\times 10^{7}$ ADD
$88.56$
$0.041\times 10^{6}$ FLOPs
$3.194\times 10^{6}$ SOPs
$3.06$ $91.62$
Ours
$7.32\times 10^{6}$SHIFT
$7.92\times 10^{6}$ADD
8.08
$0.041\times 10^{6}$ FLOPS
$2.660\times 10^{6}$ SOPs
$2.58$ $10.66$